# Conforming Spline-Based Parameterisation Techniques for Plane Graphs

### v1

@@ Abstract
This paper presents a spline-based parameterisation framework for plane graphs. The plane graph is characterised by a collection of curves forming closed loops that fence-off planar faces which have to be parameterised individually. Hereby, we focus on parameterisations that are conforming across the interfaces between the faces. Parameterising each face individually allows for the imposition of locally differing material parameters which has many applications in engineering applications, such as elasticity and heat transfer. For the parameterisation of the indidual faces, we emply the concept of harmonic maps. The plane graph’s spline-based parameterisation is suitable for numerical simulation based on isogeometric analysis or can be utilised to extract arbitrarily dense classical meshes. Application-specific features can be built into the geometry’s mathematical description either on the spline level or in the mesh extraction step.
@@ 1 Introduction
The associated difficulties are further exacerbated in settings that require a mathematical description of several geometries, in the planar case referred to as faces. Dividing a planar geometry into a set of faces is often necessary in settings that divide the geometry into regions with, for instance, differing material parameters. Since the faces may intersect at shared interfaces, most numerical techniques favour a parametric description that preserves interface conformity, i.e., wherein bordering faces’ interface elements share the same (often piecewise linear) interface curve. However, in geometrically complex settings element conformity may be difficult to accomplish which is why numerical schemes may have to be combined with mortaring techniques [1], which are themselves associated with a number of challenges.
@@ 1.1 Notation
This paper denotes vectors in boldface. The $i-th$ entry of a vector is denoted by $x_{i}$. Similarly, the $ij$-th entry of a matrix is denoted by $A_{ij}$. Let $\mathbf{y}:{\Omega}\rightarrow\mathbb{R}^{m}$ and ${\mathbf{x}}:{\Omega}\rightarrow\mathbb{R}^{n}$. Vectorial derivatives are taken along the second axis and we interchangeably employ the denotation
$\partial_{{\mathbf{x}}}\mathbf{y}:=\frac{\partial\mathbf{y}}{\partial{\mathbf{x}}}\quad\text{with}\quad\left[\frac{\partial\mathbf{y}}{\partial{\mathbf{x}}}\right]_{ij}=\frac{\partial y_{i}}{\partial x_{j}}$
Here, the vectorial derivative $\partial_{{\mathbf{x}}}\mathbf{y}$ maps into $\mathbb{R}^{m\times n}$, while $\nabla_{{\mathbf{x}}}\mathbf{y}:=(\partial_{{\mathbf{x}}}\mathbf{y})^{T}$.
Furthermore, we frequently work with vector spaces $\mathcal{V}$. By default, we employ the abuse of notation
@@
and similarly for tensorial spaces, i.e., $\mathcal{V}^{n\times n}$. Sobolev spaces are denoted by $H^{s}(\Omega)$ and vectorial Sobolev spaces by $H^{s}(\Omega,\mathbb{R}^{n})$. If a finite-dimensional space $\mathcal{V}_{h}$ is introduced via the span of its basis, we always denote the basis by $\{\mathcal{V}_{h}\}$.
@@ 1.2 Problem Statement
We are given a graph G associated with a set $V\subset\mathbb{R}^{2}$ of vertices and a set $E$ containing directed edges. Letting $\mathcal{I}_{v}=\{1,2,\ldots,N_{v}\}$ denote the index-set of vertices, we have $V=\{v_{i}\}_{i\in\mathcal{I}_{v}}$, where each $v_{i}\in V$ is a unique vertex $v_{i}\in\mathbb{R}^{2}$. The set $E=\{e_{1},\ldots,e_{N_{e}}\}$ contains edges incident to vertices. The operator $\iota:E\rightarrow\mathbb{N}\times\mathbb{N}$ returns the indices of the vertices incident to an edge $e\in E$. The set of edges connecting two vertices $\{v_{i},v_{j}\}$ is denoted by $E_{i,j}$. Typically, $|E_{i,j}|\in\{0,1\}$ but we note that $|E_{i,j}|>1$ is possible. The same operator may be applied to (ordered) sets of edges element-wise, i.e., $\iota(\{e_{1},\ldots,e_{N}\})=\{\iota(e_{1}),\ldots,\iota(e_{N})\}$.
@@
Given a planar graph $G=(V,E,\mathcal{F})$ with the aforementioned properties, the purpose of this paper is providing a framework for the spline-based parameterisation of $G$, respecting the requirement that each face $F_{i}\in\mathcal{F}$ be parameterised individually. For this purpose, we assign an $N_{F_{i}}:=|F_{i}|$-sided polygonal parametric domain ${\hat{\Omega}}_{i}\subset\mathbb{R}^{2}$ to each face $F_{i}$. By ${\hat{\Omega}}(N)$ we denote the regular $N$-sided polygon of radius one. Unless stated otherwise, we utilise ${\hat{\Omega}}_{i}={\hat{\Omega}}(N)$ with $N=|F_{i}|$.
@@ 1.5 Harmonic Maps
Finally, for $N_{\text{convex}}\geq 3$, let the corner vertices be given by $V_{\text{corner}}(F)=\left(v_{i_{1}},v_{i_{2}},v_{i_{3}},\ldots\right)\subset\mathbb{V}(F)$. The sequence $V_{\text{corner}}(F)$ naturally divides $F$ into $n\geq 3$ batches $b_{1},\ldots,b_{n}$, with $b_{j}\subset F$ comprised of the edges encountered in the sequence from $v_{i_{j}}\in V_{\text{corner}}(F)$ to $v_{i_{j+1}}\in V_{\text{corner}}(F)$ (with periodic continuation if applicable). We draw an $n$-sided polygon $\partial{\hat{\Omega}}^{\mathbf{r}}$ wherein the vertices are placed on the unit circle such that the lengths of the edges relative to $L(\partial{\hat{\Omega}}^{\mathbf{r}})$ equal the lengths of the associated batches $b_{j}$ relative to $L(\partial\Omega)$. Then, if a batch $b_{j}$ is comprised of more than one edge, the corresponding side is again segmented based on the relative lengths of the contained edges (see Figure 1, right).
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Approach 3. is a computationally inexpensive method to acquire a decent approximation of the harmonic map $\mathbf{x}^{-1}:\Omega^{S}\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ over a polygonal domain $\Omega^{S}_{h}\approx\Omega^{S}$. The approach samples ordered point sets $X=\left(x_{1},\ldots,x_{n}\right)$ and $\hat{X}=\left(\hat{x}_{1}^{r},\ldots,\hat{x}_{i}^{r}\right)$ from $\partial\Omega^{S}$ and $\partial{\hat{\Omega}}^{\mathbf{r}}$, respectively, leading to a canonical discrete correspondence $(\mathbf{f}_{h}^{\mathbf{r}\rightarrow\mathbf{x}})^{-1}:X\rightarrow\hat{X}$, with $(\mathbf{f}_{h}^{\mathbf{r}\rightarrow\mathbf{x}})^{-1}(x_{i})=\hat{x}_{i}^{r}$. The boundary of the polygon $\Omega_{h}^{S}$ is forwarded to a meshing routine which returns the triangulation $\mathcal{T}_{h}$ with boundary vertices $X\subset\mathbb{R}^{2}$ and a total number of $N>n$ vertices. An approximately harmonic map is then reconstructed from the canonical piecewise-linear $C^{0}$-continuous Lagrange basis $P_{h}(\mathcal{T}_{h})$ over $\mathcal{T}_{h}$, which leads to a linear problem of the form $Ac=Bf^{r}$. Here, the solution vector $c\in\mathbb{R}^{2(N-n)}$ contains the approximation’s weights (w.r.t. to $P_{h}(\mathcal{T}_{h})\times P_{h}(\mathcal{T}_{h})$) that are not fixed from the boundary correspondence, while the vector $f^{r}\in\mathbb{R}^{2n}$ is the concatination of the $\hat{x}_{i}^{r}\in\hat{X}$. For a computationally inexpensive method to assemble the SPD $A\in\mathbb{R}^{2N\times 2N}$ and the sparse $B\in\mathbb{R}^{2N\times 2n}$, we refer to [30]. The approximation $\left(\mathbf{x}_{h}^{\mathbf{r}}\right)^{-1}\approx\left(\mathbf{x}^{\mathbf{r}}\right)^{-1}$ maps the triangulation $\mathcal{T}_{h}$ onto a triangulation $\hat{\mathcal{T}}_{h}^{r}$ of ${\hat{\Omega}}^{\mathbf{r}}$ which automatically yields a piecewise linear approximation $\operatorname{span}P_{h}(\hat{\mathcal{T}}_{h}^{r})\times P_{h}(\hat{\mathcal{T}}_{h}^{r})\ni\mathbf{x}_{h}^{\mathbf{r}}\approx\mathbf{x}^{\mathbf{r}}$ of an inversely harmonic map from the correspondence of the vertices $\hat{\mathcal{T}}^{r}\ni\hat{v}_{i}\rightarrow v_{i}\in\mathcal{T}$. Approach 3. finds applications in Section 1.6 and Section 2.
@@
While a spline map could be fit to the vertex pairs (potentially with regularisation terms), we have found approaches 1. and 2. to be a more robust choice in practice because $\left(\mathbf{x}_{h}^{\mathbf{r}}\right)^{-1}:{\Omega}_{h}^{S}\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ may map a homogeneous triangulation $\mathcal{T}_{h}$ to a very heterogeneous triangulation $\hat{\mathcal{T}}_{h}^{r}$, which often leads to an underdetermined problem in the absence of regularisation terms.
1.6 Patch adjacency graphs
Refer to captionRefer to caption
Refer to captionRefer to caption
over all interior vertices $V^{\mathbf{r}}\ni v^{\mathbf{r}}\in V^{\mathbf{r}}_{\text{int}}$. Here, $\operatorname{ReLU}(\,\cdot\,)$ denotes the ReLU activation function $\operatorname{ReLU}(x):=\max\{0,x\}$, while $\tau\geq 0$ is an offset. We note that $C(\,\cdot\,,\tau)$ vanishes once all $\boldsymbol{\nu}(v^{\mathbf{r}},q^{\mathbf{r}})$ are above the threshold $\tau$ and the minimisation is terminated. Minimisation based on (4) is nonconvex and the cost function is nonsmooth. However, a gradient-based algorithm nevertheless converges after a small number of iterations as long as the initial guess represents a layout for which all $\boldsymbol{\nu}(v^{\mathbf{r}},q^{\mathbf{r}})$ are close to or exceed $\tau$. Starting on $\tau=0$, its value is incremented in an outer loop until optimisation no longer yields a minimiser for which $C(\mathcal{T}^{\mathbf{r}},\tau)=0$. In practice, the optimisation terminates in a fraction of a second thanks to the low number of interior vertices. Reusing the previously computed map ${\mathbf{r}}_{h}:{\hat{\Omega}}\rightarrow{\hat{\Omega}}^{\mathbf{r}}_{h}$ to create an initial layout for optimisation based on 4, in this way, hundreds of templates can be optimised in a matter of seconds.
In practice, the final step is only carried out for a single template which is selected from a large catalogue of optimised templates. The proper selection is discussed in Section 2.
All required steps are summarised by Figure 2.
@@ 2 Graph Preparation and Templatisation
We note that the main difficulty associated with the splitting of a face’s edge is the possibility of it, in turn, introducing an uneven number of edges in neighboring faces. This includes faces $F\in\mathcal{F}$ to which a suitable template $T\in\mathbb{T}$ had previously been assigned. In order to handle this situation, we introduce the concept of template refinement in Section 2.2.
@@ 2.1 Removing concave corners
To avoid concave corners, the first step of the preparation stage is generating curves originating from the concave corners that split the associated face $F_{i}\in\mathcal{F}$ into two parts $F_{i}^{+}$ and $F_{i}^{-}$, each containing only convex corners. In general, we assume that the curve connects two existing vertices $\{v_{\alpha},v_{\beta}\}\subset\mathbb{V}(F_{i})$ and is strictly contained in the interior of $\Omega_{i}$. We denote the newly-created edge between $v_{\alpha}$ and $v_{\beta}$ by $e_{\alpha\beta}$.
@@
$\left\{\begin{array}[]{l}C(0)=v_{\alpha},\\ C(1)=v_{\beta},\\ v_{\alpha}\text{ concave }\implies C^{\prime}(0)\in CC(\hat{t}_{-}(v_{\alpha}),-\hat{t}_{+}(v_{\alpha})),\\ v_{\beta}\text{ concave }\implies-C^{\prime}(1)\in CC(\hat{t}_{-}(v_{\beta}),-\hat{t}_{+}(v_{\beta})),\\ C|_{(0,1)}\subset\Omega_{i}.\end{array}\right.$ (10)
Here, $CC(v,w)$ refers to the convex cone generated by two vectors $\{v,w\}\subset\mathbb{R}^{2}$.
In the simplest case, a valid splitting curve is found by simply connecting two vertices linearly. However, we generally employ Hermite interpolation [31, Chapter 5] to generate a (piecewise) polynomial curve with specified tangent at the vertices. Given the vertex and tangent information $C(0)=v_{\alpha}$, $C^{\prime}(0)$, $C(1)=v_{\beta}$ and $C^{\prime}(1)$, the curve $C:[0,1]\rightarrow\mathbb{R}^{2}$ is given by
@@
with $\eta,\gamma>0$ and $\hat{n}(\,\cdot\,)$ as defined in Section 1.4. Heuristically, a good choice of scaling parameters is given by $\eta=\gamma=\|v_{\beta}-v_{\alpha}\|$. In practice, we flag a vertex $v_{\alpha}\in\mathbb{V}(F_{i})$ as concave whenever the discrete interior angle $\angle(v_{\alpha})$ satisfies $\angle(v_{\alpha})\geq\pi+\varepsilon_{\angle}$, where $\varepsilon_{\angle}\ll 1$ is a user-specified parameter.
In practice, a new point set $p_{\alpha\beta}:=w(e_{\alpha\beta})$ is formed by evaluating $C:[0,1]\rightarrow\overline{\Omega}_{i}$ in a large number of discrete values $\{t_{j}\}_{j}$ uniformly-spaced over the unit interval. The validity of the discrete splitting curve follows from the requirement that all interior points of $p_{\alpha\beta}$ be strictly contained in $\Omega_{i}$, which can be verified using standard polygon routines [32].
@@
$F$ $\leftarrow\left(F\setminus\{F_{i}\}\right)\cup\{F_{i}^{+},F_{i}^{-}\}.$ (13)
Letting $F_{i}=\left(e_{i_{1}},\ldots,e_{\alpha}^{-},e_{\alpha}^{+},\ldots,e_{\beta}^{-},e_{\beta}^{+},\ldots,e_{i_{N_{i}}}\right)$, where $\{e_{\alpha}^{-},e_{\alpha}^{+}\}$ and $\{e_{\beta}^{-},e_{\beta}^{+}\}$ are the two positively-oriented edges incident to $v_{\alpha}$ and $v_{\beta}$, respectively, we have
$F_{i}^{+}:=\left(e_{i_{1}},\ldots,e_{\alpha}^{-},e_{\alpha\beta},e_{\beta}^{+},\ldots,e_{i_{N_{i}}}\right)\quad\text{and}\quad F_{i}^{-}:=\left(e_{\alpha}^{+},\ldots,e_{\beta}^{-},-e_{\beta\alpha}\right).$ (14)
@@
We introduce the operator $\mathbb{V}^{\varepsilon_{\angle}}_{\text{conc}}(\,\cdot\,)$ which, when applied to an $F\in\mathcal{F}$, returns all $v_{\alpha}\in\mathbb{V}(F)$ that have been flagged concave (according to the user-specified parameter $\varepsilon_{\angle}\geq 0$). We introduce the set $F_{\text{conc}}^{\varepsilon_{\angle}}:=\{F\in\mathcal{F}\quad|\quad|\mathbb{V}^{\varepsilon_{\angle}}_{\text{conc}}(F)|>0\}$ as the set of faces containing at least one concave corner. The default strategy iterates over the $F\in F_{\text{conc}}^{\varepsilon_{\angle}}$ in an outer loop and selects pairs of vertices $\{v_{\alpha},v_{\beta}\}\subset\mathbb{V}(F)$, with |$|\{v_{\alpha},v_{\beta}\}\cap\mathbb{V}^{\varepsilon_{\angle}}_{\text{conc}}(F)|\in\{1,2\}$, as candidates for splitting. For simplicity, we restrict ourselves to a greedy iterative strategy that selects the first vertex $v_{\alpha}$ of $\mathbb{V}(F)$ in each iteration and picks the best target vertex $v_{\beta}\in\mathbb{V}(F)\setminus\{v_{\alpha}\}$ as measured by a quality criterion. Given some $v_{\alpha}\in\mathbb{V}^{\varepsilon_{\angle}}_{\text{conc}}(F)$, the criterion first filters all pairs that lead to a valid point set $p_{\alpha\beta}$ when sampling a cubic Hermite curve that follows from (11) and (12) in a large number of uniformly-spaced points $\{t_{j}\}_{j}$ over the unit interval. Here, we always impose the tangent information $C^{\prime}(0)=-\eta\hat{n}(v_{\alpha})$ and $C^{\prime}(1)=\gamma\hat{n}(v_{\beta})$ even when $v_{\beta}$ is not a concave vertex. Then, a measure of quality is assigned to all remaining pairs. Heuristically, the best results are obtained for curves $C_{\alpha\beta}$ that minimise a measure of their deviation from a straight line. As such, we measure the quality of the splitting between $v_{\alpha}$ and $v_{\beta}$ by the cost function
@@
Heuristically, a good choice is $\mu=2$. The greedy approach for the removal of concave corners is given in Algorithm 1. It takes as input the graph $G=(V,E,\mathcal{F})$, the concavity threshold $\varepsilon_{\angle}\geq 0$ and the parameter $\mu\geq 1$ we introduced in (18). Then it attempts to remove all concave corners and returns the indices of the faces that contain at least one concave corner for which no suitable splitting curve was found. Since the faces are treated independently of one another, the order in which the algorithm iterates over concave faces is unimportant.

### v2

# On the Spline-Based Parameterisation of Plane Graphs via Harmonic Maps

@@ Abstract
This paper presents a spline-based parameterisation framework for plane graphs. The plane graph is characterised by a collection of curves forming closed loops that fence-off planar faces which have to be parameterised individually. Hereby, we focus on parameterisations that are conforming across the interfaces between the faces. Parameterising each face individually allows for the imposition of locally differing material parameters which has applications in various engineering disciplines, such as elasticity and heat transfer. For the parameterisation of the individual faces, we employ the concept of harmonic maps. The plane graph’s spline-based parameterisation is suitable for numerical simulation based on isogeometric analysis or can be utilised to extract arbitrarily dense classical meshes. Application-specific features can be built into the geometry’s mathematical description either on the spline level or in the mesh extraction step.
@@ 1 Introduction
The associated difficulties are further exacerbated in settings that require a mathematical description of several geometries, in the planar case referred to as faces. Dividing a planar geometry into a set of faces is necessary in settings that impose, for instance, locally differing material parameters. Since the faces may intersect at shared interfaces, most numerical techniques favour a parametric description that preserves interface conformity. That is, wherein bordering faces’ interface elements share the same (often piecewise linear) interface curve. However, in geometrically complex settings element conformity may be difficult to accomplish which is why numerical schemes may have to be combined with mortaring techniques [2], which are themselves associated with a number of challenges.
@@ 1.1 Notation
This paper denotes vectors in boldface. The $i-th$ entry of a vector is denoted by $x_{i}$. Similarly, the $ij$-th entry of a matrix is denoted by $A_{ij}$. Let $\mathbf{y}:{\Omega}\rightarrow\mathbb{R}^{m}$ and ${\mathbf{x}}:{\Omega}\rightarrow\mathbb{R}^{n}$. We frequently work with vector spaces $\mathcal{V}$. By default, we employ the abuse of notation
@@
and similarly for tensorial spaces, i.e., $\mathcal{V}^{n\times n}$. Sobolev spaces are denoted by $H^{s}(\Omega)$ and vectorial Sobolev spaces by $H^{s}(\Omega,\mathbb{R}^{n})$.
@@ 1.2 Problem Statement
We are given a graph G associated with a set $V\subset\mathbb{R}^{2}$ of vertices and a set $E$ containing directed edges. Letting $\mathcal{I}_{v}=\{1,2,\ldots,N_{v}\}$ denote the index-set of vertices, we have $V=\{v_{i}\}_{i\in\mathcal{I}_{v}}$, where each $v_{i}\in V$ is a unique vertex $v_{i}\in\mathbb{R}^{2}$. The set $E=\{e_{1},\ldots,e_{N_{e}}\}$ contains edges incident to vertices. The operator $\iota:E\rightarrow\mathbb{N}\times\mathbb{N}$ returns the indices of the vertices incident to an edge $e\in E$. The set of edges connecting two vertices $\{v_{i},v_{j}\}$ is denoted by $E_{i,j}$. Typically, $|E_{i,j}|\in\{0,1\}$ but we note that $|E_{i,j}|>1$ is possible, see Figure 1. The same operator may be applied to (ordered) sets of edges element-wise, i.e., $\iota(\{e_{1},\ldots,e_{N}\})=\{\iota(e_{1}),\ldots,\iota(e_{N})\}$.
@@
Figure 1: An example of a plane graph (left) that is visualised via its weight function $w(\,\cdot\,)$. The various polygons $\Omega_{i}$ that are bounded by the point sets $w(e),\,e\in F_{i}$ are highlighted in different colors. The graph contains several examples of two edges connecting the same two vertices, for example $e_{10}$ and $e_{13}$. The figure shows that $F_{1}=\left(\pm e_{1},\pm e_{2},\ldots,\pm e_{12}\right)$. The right figure shows a possible spline-based parameterisation wherein each point set $w(e)$ is replaced by an accurate spline approximation $w^{S}(e)$ fencing off the spline domains $\Omega_{i}^{S}\approx\Omega_{i}$. Each face is parameterised from one or several patches using a harmonic map, while neighboring faces use the same spline edges $w^{S}(e)$, thus leading to a conforming interface.
@@
Given a planar graph $G=(V,E,\mathcal{F})$ with the aforementioned properties, the purpose of this paper is providing a framework for the spline-based parameterisation of $G$, respecting the requirement that each face $F_{i}\in\mathcal{F}$ be parameterised individually, see Figure 1 right. For this purpose, we assign an $N_{F_{i}}:=|F_{i}|$-sided polygonal parametric domain ${\hat{\Omega}}_{i}\subset\mathbb{R}^{2}$ to each face $F_{i}$. By ${\hat{\Omega}}(N)$ we denote the regular $N$-sided polygon of radius one. Unless stated otherwise, we utilise ${\hat{\Omega}}_{i}={\hat{\Omega}}(N)$ with $N=|F_{i}|$.
@@ 1.5 Harmonic Maps
Finally, for $N_{\text{convex}}\geq 3$, let the corner vertices be given by $V_{\text{corner}}(F)=\left(v_{i_{1}},v_{i_{2}},v_{i_{3}},\ldots\right)\subset\mathbb{V}(F)$. The sequence $V_{\text{corner}}(F)$ naturally divides $F$ into $n\geq 3$ batches $b_{1},\ldots,b_{n}$, with $b_{j}\subset F$ comprised of the edges encountered in the sequence from $v_{i_{j}}\in V_{\text{corner}}(F)$ to $v_{i_{j+1}}\in V_{\text{corner}}(F)$ (with periodic continuation if applicable). We draw an $n$-sided polygon $\partial{\hat{\Omega}}^{\mathbf{r}}$ wherein the vertices are placed on the unit circle such that the lengths of the edges relative to $L(\partial{\hat{\Omega}}^{\mathbf{r}})$ equal the lengths of the associated batches $b_{j}$ relative to $L(\partial\Omega)$. Then, if a batch $b_{j}$ is comprised of more than one edge, the corresponding side is again segmented based on the relative lengths of the contained edges (see Figure 2, right). This ensures that the induced correspondence $\mathbf{f}^{{\mathbf{r}}\rightarrow{\mathbf{x}}}:\partial{\hat{\Omega}}^{\mathbf{r}}\rightarrow\partial{\Omega}^{S}$, with $\partial{\Omega}^{S}\approx\partial{\Omega}$ the spline approximation, has a nearly continuous tangent on the straight segments of $\partial{\Omega}^{\mathbf{r}}$.
@@
Contrary to approaches 1. and 2., approach 3. does not operate on the spline input but on discrete point cloud approximations $(\partial{\hat{\Omega}}^{\mathbf{r}}_{h},\partial\Omega_{h}^{S})$ sampled from $(\partial{\hat{\Omega}}^{\mathbf{r}},\partial\Omega^{S})$. The algorithm then utilises a triangulation of $\Omega^{S}_{h}$ to provide a piecewise-linear map $\mathbf{x}_{h}^{\mathbf{r}}:{\hat{\Omega}}^{\mathbf{r}}_{h}\rightarrow{\Omega}^{S}_{h}$ whose inverse is approximately harmonic. Approach 3. leads to a linear problem that is computationally inexpensive compared to approaches 1. and 2. It finds applications as a surrogate for $\mathbf{x}^{\mathbf{r}}:{\hat{\Omega}}^{\mathbf{r}}\rightarrow\Omega^{S}$ in Section 2 and is discussed in greater detail in A.
@@
While a spline map could be fit to the vertex pairs (potentially with regularisation terms) created by approach 3., we have found approaches 1. and 2. to be a more robust choice in practice because $\left(\mathbf{x}_{h}^{\mathbf{r}}\right)^{-1}:{\Omega}_{h}^{S}\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ may map a homogeneous triangulation $\mathcal{T}_{h}$ to a very heterogeneous triangulation $\hat{\mathcal{T}}_{h}^{r}$, which often leads to an underdetermined problem in the absence of regularisation terms.
1.6 Creation of a Template Catalogue
Approaches 1. and 2. from Section 1.5 operate on a multipatch spline space $\mathcal{V}_{h}\subset H^{1}({\hat{\Omega}}_{i})$ defined over a quadrangulation of ${\hat{\Omega}}_{i}$ which is itself represented by a template $\mathbb{T}\ni T_{i}=(V_{i},E_{i},\mathcal{Q}_{i})$ with four-sided faces $q\in\mathcal{Q}_{i}$, see Figure 3.
Figure 3: Figure showing a template $\mathbb{T}\ni T_{i}=(V_{i},E_{i},\mathcal{Q}_{i})$ for the $8$-sided regular polygon of radius one. The figure additionally shows the corresponding multipatch covering of a lens-shaped control domain ${\hat{\Omega}}^{\mathbf{r}}$ (see Section 1.5). The control domain is parameterised by the controlmap $\mathbf{r}:{\hat{\Omega}}(8)\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ that is constructed using optimisation of the inner patch vertices along with the Coons’ patch approach. For details, see B.
Using the concept of patch adjacency graphs (PAGs), as introduced in [4], for this paper we generated a template catalogue $\mathcal{T}$ comprised of $\approx 6\times 10^{4}$ templates covering the regular $N$-sided polygon ${\hat{\Omega}}(N)$ of radius one for $N\in\{4,6,\ldots,16\}$. We note that the computation of all PAGs along with the generation of a concrete template $T_{i}=(V_{i},E_{i},\mathcal{Q}_{i})$ from the PAG is an offline computation which does not contribute to the computational costs of what follows.
Since each domain $\Omega$ that follows from some $F\in\mathcal{F}$ is accompanied by a canonical control domain ${\hat{\Omega}}^{\mathbf{r}}$ (see Section 1.5), a template $T_{i}\in\mathbb{T}$ covering ${\hat{\Omega}}(N)$ additionally requires the creation of a piecewise smooth (over the quadrangular faces) control map $\mathbf{r}:{\hat{\Omega}}(N)\rightarrow{\hat{\Omega}}^{\mathbf{r}}$. Unlike the creation of $\mathbb{T}$ this step constitutes an online computation and is tackled via the computationally inexpensive minimisation of a functional.
The concept of PAGs, the creation of the catalogue $\mathbb{T}$ and the computation of a valid controlmap $\mathbf{r}:{\hat{\Omega}}(N)\rightarrow{\Omega}^{\mathbf{r}}$ for layouts $T_{i}\in\mathbb{T}$ is discussed in B.
@@ 2 Graph Preparation and Templatisation
We note that the main difficulty associated with the splitting of a face’s edge is the possibility of it, in turn, introducing an uneven number of edges in a neighbouring face. This includes faces $F\in\mathcal{F}$ to which a suitable template $T\in\mathbb{T}$ had previously been assigned. In order to handle this situation, we introduce the concept of template refinement in Section 2.2.
@@ 2.1 Removing concave corners
To avoid concave corners, the first step of the preparation stage is repeatedly generating curves that originate in a concave vertex $v\in\mathbb{V}(F),F\in\mathcal{F}$ while splitting $F$ into two parts $F^{+}$ and $F^{-}$ that no longer contain the concave $v\in\mathbb{V}(F)$. In general, we assume that the curve connects two existing vertices $\{v_{\alpha},v_{\beta}\}\subset\mathbb{V}(F_{i})$ and is strictly contained in the interior of $\partial{\Omega}=\cup w(e),e\in F$. We denote the newly-created edge between $v_{\alpha}$ and $v_{\beta}$ by $e_{\alpha\beta}$.
@@
$\left\{\begin{array}[]{l}C(0)=v_{\alpha},\\ C(1)=v_{\beta},\\ v_{\alpha}\text{ concave }\implies C^{\prime}(0)\in CC(\hat{t}_{-}(v_{\alpha}),-\hat{t}_{+}(v_{\alpha})),\\ v_{\beta}\text{ concave }\implies-C^{\prime}(1)\in CC(\hat{t}_{-}(v_{\beta}),-\hat{t}_{+}(v_{\beta})),\\ C|_{(0,1)}\subset\Omega.\end{array}\right.$ (8)
Here, $CC(v,w)$ refers to the convex cone generated by two vectors $\{v,w\}\subset\mathbb{R}^{2}$, see Figure 4.
Figure 4: An example of removing a concave vertex (blue) via the use of Hermite splitting curves whose starting and end point tangents are contained in the endpoints’ convex cones. The figure shows all possible Hermite curves originating in the bottom concave vertex $v\in V$. While all curves are valid, selection based on (16) would favour the curve highlighted in blue since it is the straightest and removes two concave vertices at once.
In the simplest case, a valid splitting curve is found by simply connecting two vertices linearly. However, we generally employ Hermite interpolation [20, Chapter 5] to generate a (piecewise) polynomial curve with specified tangent at the vertices. Given the vertex and tangent information $C(0)=v_{\alpha}$, $C^{\prime}(0)$, $C(1)=v_{\beta}$ and $C^{\prime}(1)$, the curve $C:[0,1]\rightarrow\mathbb{R}^{2}$ is given by
@@
with $\eta,\gamma>0$ and $\hat{n}(\,\cdot\,)$ as defined in Section 1.4. Heuristically, a good choice of scaling parameters is given by $\eta=\gamma=\|v_{\beta}-v_{\alpha}\|$. We flag a vertex $v_{\alpha}\in\mathbb{V}(F)$ as concave whenever the discrete interior angle $\angle(v_{\alpha})$ satisfies $\angle(v_{\alpha})\geq\pi+\varepsilon_{\angle}$, where $\varepsilon_{\angle}\ll 1$ is a user-specified parameter.
In practice, a new point set $p_{\alpha\beta}:=w(e_{\alpha\beta})$ is formed by evaluating $C:[0,1]\rightarrow\overline{\Omega}$ in a large number of discrete values $\{t_{j}\}_{j}$ uniformly-spaced over the unit interval. The validity of the discrete splitting curve follows from the requirement that all interior points of $p_{\alpha\beta}$ be strictly contained in $\Omega$, which can be verified using standard polygon routines [27].
@@
$\mathcal{F}$ ($\leftarrow\left(\mathcal{F}\setminus\{F\}\right)\cup\{F^{+},F^{-}\}.$ (11)
Letting $F=\left(e_{i_{1}},\ldots,e_{\alpha}^{-},e_{\alpha}^{+},\ldots,e_{\beta}^{-},e_{\beta}^{+},\ldots,e_{i_{N_{i}}}\right)$, where $\{e_{\alpha}^{-},e_{\alpha}^{+}\}$ and $\{e_{\beta}^{-},e_{\beta}^{+}\}$ are the two positively-oriented edges incident to $v_{\alpha}$ and $v_{\beta}$, respectively, we have
$F^{+}:=\left(e_{i_{1}},\ldots,e_{\alpha}^{-},e_{\alpha\beta},e_{\beta}^{+},\ldots,e_{i_{N_{i}}}\right)\quad\text{and}\quad F^{-}:=\left(e_{\alpha}^{+},\ldots,e_{\beta}^{-},-e_{\beta\alpha}\right).$ (12)
@@
We introduce the operator $\mathbb{V}^{\varepsilon_{\angle}}_{\text{conc}}(\,\cdot\,)$ which, when applied to an $F\in\mathcal{F}$, returns all $v_{\alpha}\in\mathbb{V}(F)$ that have been flagged concave (according to the user-specified parameter $\varepsilon_{\angle}\geq 0$). We introduce the set $\mathcal{F}_{\text{conc}}^{\varepsilon_{\angle}}:=\{F\in\mathcal{F}\quad|\quad|\mathbb{V}^{\varepsilon_{\angle}}_{\text{conc}}(F)|>0\}$ as the set of faces containing at least one concave corner. The default strategy is based on a greedy algorithm that iterates over the $F\in\mathcal{F}_{\text{conc}}^{\varepsilon_{\angle}}$ in an outer loop and selects pairs of vertices $\{v_{\alpha},v_{\beta}\}\subset\mathbb{V}(F)$, with |$|\{v_{\alpha},v_{\beta}\}\cap\mathbb{V}^{\varepsilon_{\angle}}_{\text{conc}}(F)|\in\{1,2\}$, as candidates for splitting.
Given some $F\in\mathcal{F}^{\varepsilon_{\angle}}_{\text{conc}}$, let $E_{\text{valid}}$ be the set of edges $e_{\alpha\beta}$ between $\{v_{\alpha},v_{\beta}\}\subset\mathbb{V}(F)$, with |$|\{v_{\alpha},v_{\beta}\}\cap\mathbb{V}^{\varepsilon_{\angle}}_{\text{conc}}(F)|\in\{1,2\}$ that result in a valid splitting curve $C_{\alpha\beta}$ (sampled from (9)). We measure the quality of $C_{\alpha\beta}$ using the cost function
@@
Heuristically, a good choice is $\mu=2$. The greedy strategy now splits $F\in\mathcal{F}^{\varepsilon_{\angle}}_{\text{conc}}$ using the edge $e_{\alpha\beta}\in E_{\text{valid}}$ that minimises (16). The graph and weight function $w(\,\cdot\,)$ are updated to reflect the changes and the greedy strategy is called on the updated graph until no more concave vertices that are eligible for splitting are found. Should algorithm fail to remove a concave corner, a valid splitting curve is created manually.
The greedy strategy is summarised in Algorithm 1.